# Emerging Technologies for 6G Non-Terrestrial-Networks: From Academia to Industrial Applications

### v1

@@ III NETWORK SLICING-AIDED NTN
Network slicing is a powerful technique in telecommunications, which allows operators to create multiple virtual networks from a shared physical network architecture. Utilizing technologies such as SDN and NFV, the operators can dynamically allocate network resources to each network slice according to its specific requirements in terms of bandwidth, latency, and reliability. In the context of NTN, network slicing enables a single network infrastructure to serves multiple types of applications and users with different demands. For example, remote control for UAVs often requires low-latency communication, whereas satellite internet service might require higher bandwidth to serve millions of users. For both cases, network slicing can divide a single physical NTN into two virtual slices to meet different demands by allocating different resources to each slice.
III-A Satellite Networks
@@ III-A1 Resource Allocation
Among those problems, resource-allocation-related challenges have been addressed in [49, 50, 51, 52]. Specifically, in [49], the authors propose innovative approaches to integrate network slicing into railway communication. These approaches aim to optimize network resource allocation by ensuring that latency requirements are met for each class of service in a network. First, a mathematical approach, namely Queuing Theory (QT) method, uses a nonlinear constrained optimization to minimize the difference between the latency. In addition, a Neural Network (NN) approach, namely NN total model, utilizes delays obtained through simulations as input dataset to produce the control variables. Since this approach approximates the required latency by simple fitting, it could lead to inefficient and inaccurate results. To overcome this drawback, the authors propose an NN-divided model [49] to approximate the latency of satellite and terrestrial networks using two separate neural networks. This approach solves an optimization problem similar to that of the QT-based method, except that it minimizes an expected mean latency. Simulation results show that these approaches can reduce the latency by roughly 25%-60%. With a focus on dynamic allocation of radio resources, in [50], the authors propose a RL-based slicing strategy, namely dynamic radio resource slicing strategy (Sat-RRSlice), to serve the LEO networks. Particularly, the network slicing problem is modeled as a Semi-Markov Decision Process (SMDP), including network states, actions, state dynamics, and rewards. The states of each slice denote the resource allocation and utilization at a time slot. Moreover, the actions, taken by a resource manager on each slice at a time slot, indicate whether a resource unit is allocated to a slice or not. Furthermore, the reward function is used to calculate the reward, defined by the network’s total utility, for taking a certain action. Using Q-learning, optimal Q-values are learned iteratively based on information obtained from a wireless environment over discrete time slots. Simulation results show that, compared with the static slicing strategy [50], Sat-RRSlice can increase the resource utilization rate and sum utility (i.e., the total benefit gained from communication) by at least 18.5% and 9%, respectively.
@@
Taking another approach, in [51], the authors design a network-slicing architecture for LEO satellite network with a focus on increasing flexibility and resource utilization. In this architecture, network slicing is managed through Network Slice Selection Assistance Information (NSSAI) and Network Slice Selection Policy (NSSP). The NSSAI, stored in each device, is specifically configured for each device while the NSSP is the set of rules that the network management component must follow. Moreover, two types of network slices are defined: shared slices used by all users for common functions (e.g., storing user data and authenticating users), and dedicated slices for providing personalized service. Experimental results show that a network implemented with this architecture can effectively handle multimedia communications over IP networks. The authors in [52] design a framework for SDN/NFV-enabled satellite ground segment systems to enable on-demand network slicing. This framework, namely OnDReAMS, has a Service Orchestrator (SO) responsible for managing the life cycle of network slices, e.g., instantiation, maintenance, and termination. Besides, an NFV Manager is used to handle the instantiation, modification, and termination of the VNFs. Moreover, a novel component, namely Satellite Network Slice Descriptor (SNSD), describes the characteristics of the slice as requested by the customer, allowing the SO to set up the slice with flexibility. For resource allocation, the slicing problem is solved using a Mixed Integer Linear Program (MILP) model. A time-window-based online algorithm is then developed to handle the on-demand aspect of the solution, which solves the MILP at the beginning of each time window with updated slice resources. Simulation results show that OnDReAMS can reduce the average of QoS violation by approximately 10%-50% compared with a baseline approach.
III-A2 Routing
@@
III-A3 Satellite Edge Computing
Network slicing application to satellite edge computing (SatEC) architectures are discussed in [47] and [46]. Particularly, in [47], the authors present an IoT-supportable SatEC architecture to use satellites for 6G IoT services efficiently. To that end, the authors propose solutions to address two problems: balance the trade-offs between latency and power, and managing network resource allocation. For the first problem, a multi-objective optimization problem is formulated considering latency, computational power, and transmission power attenuation. A Satellite Edge Multi-objective Tabu Search (SE-MOTS) [55] algorithm is used to find a Pareto-optimal point, which presents the best trade-off among the three factors considering the dynamic satellite topology and service requirements. For the second aspect, a sliced SatEC optimization problem, formulated as a normalized weighted sum of three objective functions, is used to schedule the tasks with different demands. A golden-section method [56] is used to solve this aspect with low computational resources. Taking another approach, in [46], the authors design a system architecture that combines space-based edge computing and network slicing for space-based network resource management, as illustrated in Fig. 9. In this architecture, edge computing nodes are included due to their benefits of being closer to the data sources and their ability to process data locally. These nodes, namely distributed fog satellite nodes and centralized space-based edge clouds, are responsible for computing tasks with low and high complexity, respectively. In the control plane of the network, the authors introduce the two-layer controller, Fog SDN controller (FSC) and Cloud SDN controller & Slicing Manager (CSC&SM). Particularly, FSC abstracts fog satellite nodes and manages their resources while CSC&SM optimizes resource management of network slices with a global view of the network. Moreover, a 5G Satellite Network Slice Management(5G SNSM) [46] is proposed to perform flexible nerwork slice allocation based on QoS requirements. Simulation results show that the proposed architecture can improve the end-to-end delay and jitter by at least 7% and 50%, respectively.
@@ III-A3 Satellite Edge Computing
III-A4 Slice Admission Control
Slice admission control is another important aspect of network slicing that should be considered simultaneously with resource allocation. In [57], a two-stage approach is developed to jointly optimize resource allocation and slice admission control in LEO satellite networks. To this end, the authors first formulate a robust optimization problem to optimize slice admission, taking into account the uncertainty in the network state and user distribution. Based on the admission decisions in the first stage, the resource allocation strategies are optimized in the second stage to guarantee QoS for users. Since the original optimization problem is non-convex, auxiliary variables are introduced to transform the problem into a convex form, thereby significantly reducing the problem’s complexity. Simulations results show that the proposed approach can improve the user QoS by 7.2% compared to the non-robust approach.
III-B Integrated Satellite-Terrestrial Networks
@@ III-B1 Resource Allocation
Innovative approaches are proposed to overcome those challenges. To address the resource allocation problem in slice-aware ISTN, optimizing the VNE [60] algorithm is a promising solution. To that end, the authors in [59] design a framework, namely Slice-Aware VNE for Satellite-Terrestrial (SAST-VNE) to improve the implementation of the VNE algorithm in ISTNs. In this framework, problematic network slices, i.e., those that do not meet Key Performance Indicators (KPIs) or will undergo satellite handover, are inserted into a queue for analysis. In case a new slice is created, the embedding is computed using the DViNE algorithm [61]. For problematic network slices, if they have high priority, i.e., low tolerated latency, each affected link will be remapped with the shortest-path algorithm. On the contrary, if they have lower priority, the authors propose an interative algorithm, namely SAST-VNE, to balance the network load and minimize the migration cost. Simulation results show that during handovers, SAST-VNE can reduce the average node migrations and average link migration by roughly 25% and 6%, respectively. Another implementation of VNE for ISTNs is presented in [58]. Particularly, the authors design a testbed to validate the feasibility of integrating non-GEO satellite constellations and implementing VNE algorithms in highly dynamic network conditions. This testbed consists of a dynamic satellite-terrestrial network emulated in Mininet [58], an external Ryu SDN [58] controller and a VNE algorithm script. First, the Mininet emulator is used to build an OpenFlow-based substrate network, enabling the Ryu SDN controller to manage the flow of packets in the network. The Ryu SDN controller also uses a Traffic Engineering (TE) application to create paths for each VN, set rate limits, gather network statistics, read the network topology, and process changes to the network topology in real-time. Furthermore, the VNE algorithm script, implemented in Matlab using an Integer Linear Programming formulation with a load-balancing objective function, handles required resources for virtual networks and updates the network configuration to the Ryu SDN controller. Simulation results show that the testbed can handle many scenarios and changes in the network topology such as adding new VNs, dynamic changes in the network layout, handling of network failures.
@@
Moreover, approaches to optimize resource allocation are presented in [62] and [63]. Particularly, the authors in [62] propose a dynamic slicing strategy for ISTNs. The proposed strategy uses mirror nodes, instead of virtual nodes, to store motion duration and resource information of satellites. Based on that, service forwarding path and slice resources can be adjusted, as illustrated in Fig. 10. The mirror nodes also communicate the satellite resources with ground stations and receive feedback, i.e., which satellites to be chosen for the slicing process. An optimization model is then formulated to maximize the long-term average slicing performance of the network with a constraint of computing, storage and bandwidth resources. Moreover, a resource reserved in slices and adaptive adjustment between slices algorithm (R2A2) is developed to solve the optimization problem. Simulation results show that, compared to a DQN approach, R2A2 can increase resource utilization and slice satisfaction by roughly 33% and 30%, respectively. With a focus on admission decision for a slice request, the authors in [63] propose a game-theory-based solution for resource allocation in ISTNs. To solve the challenges related to limited resource constraints, slice admission control is formulated involving different combinations of service providers and users. Specifically, service providers and users form a multi-sided market to exchange resources. To participate in the network, both sides consider their prices, i.e., network resources for the providers and the cost for the users. An auction, designed with the Multi-Sided Ascending-Price Auction Mechanism [63], is performed by increasing the price for both sides based on supply and demand until a balance is reached. Eventually, the final prices are set, and the allocation of resources to users is determined. Experimental results show that, compared to a baseline approach, i.e., second-price auction [63], the ascending-price auction can increase the bandwidth per user, the admission ratio, and the gain from trade per user by up to 25%, 33%, and 20%, respectively.
@@
A novel approach for network slice orchestration is presented in [64]. Particularly, the authors present a system model for deployment of network slicing for aircrafts using satellites. This model aims to enhance resource management in ISTNs by incorporating the concept of slice collaboration, i.e., defining network slices based on their willingness to share user traffic statistics with the infrastructure provider). To this end, a Mixed Integer Non Linear Programming (MINLP) model is developed to maximize the resources, i.e., cache and backhaul resources) while satisfying slice constraints. Moreover, a pricing model [64] is proposed to increase the chances of uncooperative slices, i.e., slices that are unwilling to share user traffic statistics, being served. Simulation results show that the MINLP model can increase the selection probability of uncooperative slices by 33% and cooperative slicing can accommodate 200% more services compared to that of uncooperative slicing. Taking another approach, the authors in [28] propose a slice-aware NTN architecture to enable the seamless integration between NTN and terrestrial networks. To this end, the authors propose an end-to-end slicing model which treats the NTN as a slice-aware link in the terrestrial network. Based on the proposed model, a functional architecture fully compliant with the 3GPP standard is developed to allow the NTN to be integrated with the terrestrial network. The architecture consists of three main segments: the user segment with the satellite terminal, the ground segment with satellite gateways and network control components, and the space segment with one or more satellites. The Satellite Terminal (ST) serves as the User Equipment (UE) in 5G networks, communicating with the space segment, which then relays messages to the satellite gateway. Satellite gateways are the central entities that manage resource allocation, authentication, and data processing. Simulation results show that the proposed slice-aware scheme can achieve better trip time, packet error rate, and jitter compared to those of the cases without network slicing.
III-B2 User Assignment
@@
In addition to routing, several frameworks have been proposed to enhance network performance by optimizing user assignment in [65], [41], and [66]. Particularly, in [65], the authors propose a network slicing scheme for hybrid satellite-terrestrial networks, aiming to improve reliability and reduce the video traffic offload. To this end, the authors develop a scheduling strategy that assign the users to three different types of networks, e.g., only satellite, only terrestrial, and hybrid satellite-terrestrial, based on Channel Quality Indicator (CQI) and QoS requirements. Simulation results show that the proposed scheduling strategy can improve the network throughput by up to 48% compared to the case where only the satellite network is employed. Differently, with a focus on space-terrestrial integrated vehicular networks (STIVN), the authors in [41] propose a scheme for network slicing to support both delay-tolerant services (DTSs) and delay-sensitive services (DSSs). This scheme aims to solve the problem of scheduling and resource slicing in STIVN, which involves allocating spectrum resources to slices, determining bandwidth allocation, and user assignment for each vehicle. Specifically, the system cost (i.e., DSS requirement violation, DTS delay, and slice reconfiguration) is taken into account in two subproblems: a resource slicing subproblem in large-timescale, and a resource scheduling subproblem in smaller timescale. To solve those subproblems, a two-layered RL-based scheme is developed. In the first layer, namely resource slicing layer, spectrum-resources are pre-allocated to slices using a proximal policy optimization (PPO)-based RL algorithm [41]. In the second layer, namely resource scheduling layer, depending on network conditions and service requirements, spectrum resources are assigned to vehicles using match-based algorithms. Simulation results show that the proposed approach can reduce the overall system cost up to roughly 72.57%, compared to that of a baseline approach. In another approach, the authors in [66] propose a network slicing scheme to improve their existing resource management approach in satellite-LTE networks. In the previous approach, an adaptive hybrid satellite-LTE downlink scheduler (H-MUDoS) [67] determines if users can be served through satellite network or ground-based stations. However, simulation results show that the decrease in QoS of the satellite network affects the performance of the entire hybrid network. In the current approach, the network is separated into isolation slices, each with its own scheduling strategy assigned by the scheduler. Simulation results show that the decrease in QoS of a slice does not affect other slices.
III-B3 Routing

### v2

@@ 3 Network Slicing-aided NTN
Network slicing is a powerful technique in telecommunications, which allows operators to create multiple virtual networks from a shared physical network architecture. Utilizing technologies such as SDN and NFV, the operators can dynamically allocate network resources to each network slice according to its specific requirements in terms of bandwidth, latency, and reliability. In the context of NTN, network slicing enables a single network infrastructure to serve multiple types of applications and users with different demands. For example, remote control for UAVs often requires low-latency communication, whereas satellite internet service might require higher bandwidth to serve millions of users. For both cases, network slicing can divide a single physical NTN into two virtual slices to meet different demands by allocating different resources to each slice.
3.1 Satellite Networks
@@ 3.1.1 Resource Allocation
Among those problems, resource-allocation-related challenges have been addressed in [71, 72, 73, 74]. Specifically, in [71], the authors propose innovative approaches to integrate network slicing into railway communication. These approaches aim to optimize network resource allocation by ensuring that latency requirements are met for each class of service in a network. First, a mathematical approach, namely Queuing Theory (QT) method, uses a nonlinear constrained optimization to minimize the difference between the latency. In addition, a Neural Network (NN) approach, namely NN total model, utilizes delays obtained through simulations as input dataset to produce the control variables. Since this approach approximates the required latency by simple fitting, it could lead to inefficient and inaccurate results. To overcome this drawback, the authors propose an NN-divided model [71] to approximate the latency of satellite and terrestrial networks using two separate neural networks. This approach solves an optimization problem similar to that of the QT-based method, except that it minimizes an expected mean latency. Simulation results show that these approaches can reduce the latency by roughly 25%-60%. With a focus on the dynamic allocation of radio resources, in [72], the authors propose an RL-based slicing strategy, namely dynamic radio resource slicing strategy (Sat-RRSlice), to serve the LEO networks. Particularly, the network slicing problem is modeled as a Semi-Markov Decision Process (SMDP), including network states, actions, state dynamics, and rewards. The states of each slice denote the resource allocation and utilization at a time slot. Moreover, the actions, taken by a resource manager on each slice at a time slot, indicate whether a resource unit is allocated to a slice or not. Furthermore, the reward function is used to calculate the reward, defined by the network’s total utility, for taking a certain action. Using Q-learning, optimal Q-values are learned iteratively based on information obtained from a wireless environment over discrete time slots. Simulation results show that, compared with the static slicing strategy [72], Sat-RRSlice can increase the resource utilization rate and sum utility (i.e., the total benefit gained from communication) by at least 18.5% and 9%, respectively.
@@
Taking another approach, in [73], the authors design a network-slicing architecture for the LEO satellite network with a focus on increasing flexibility and resource utilization. In this architecture, network slicing is managed through Network Slice Selection Assistance Information (NSSAI) and Network Slice Selection Policy (NSSP). The NSSAI, stored in each device, is specifically configured for each device while the NSSP is the set of rules that the network management component must follow. Moreover, two types of network slices are defined: shared slices used by all users for common functions (e.g., storing user data and authenticating users), and dedicated slices for providing personalized service. Experimental results show that a network implemented with this architecture can effectively handle multimedia communications over IP networks. The authors in [74] design a framework for SDN/NFV-enabled satellite ground segment systems to enable on-demand network slicing. This framework, namely OnDReAMS, has a Service Orchestrator (SO) responsible for managing the life cycle of network slices, e.g., instantiation, maintenance, and termination. Besides, an NFV Manager is used to handle the instantiation, modification, and termination of the VNFs. Moreover, a novel component, namely Satellite Network Slice Descriptor (SNSD), describes the characteristics of the slice as requested by the customer, allowing the SO to set up the slice with flexibility. For resource allocation, the slicing problem is solved using a Mixed Integer Linear Program (MILP) model. A time-window-based online algorithm is then developed to handle the on-demand aspect of the solution, which solves the MILP at the beginning of each time window with updated slice resources. Simulation results show that OnDReAMS can reduce the average QoS violation by approximately 10%-50% compared with a baseline approach.
3.1.2 Routing
@@
3.1.3 Satellite Edge Computing
Network slicing applications to satellite edge computing (SatEC) architectures are discussed in [69] and [68]. Particularly, in [69], the authors present an IoT-supportable SatEC architecture to use satellites for 6G IoT services efficiently. To that end, the authors propose solutions to address two problems: balancing the trade-offs between latency and power and managing network resource allocation. For the first problem, a multi-objective optimization problem is formulated considering latency, computational power, and transmission power attenuation. A Satellite Edge Multi-objective Tabu Search (SE-MOTS) [77] algorithm is used to find a Pareto-optimal point, which presents the best trade-off among the three factors considering the dynamic satellite topology and service requirements. For the second aspect, a sliced SatEC optimization problem, formulated as a normalized weighted sum of three objective functions, is used to schedule the tasks with different demands. A golden-section method [78] is used to solve this aspect with low computational resources. Taking another approach, in [68], the authors design a system architecture that combines space-based edge computing and network slicing for space-based network resource management, as illustrated in Fig. 9. In this architecture, edge computing nodes are included due to their benefits of being closer to the data sources and their ability to process data locally. These nodes, namely distributed fog satellite nodes and centralized space-based edge clouds, are responsible for computing tasks with low and high complexity, respectively. In the control plane of the network, the authors introduce the two-layer controller, Fog SDN controller (FSC) and Cloud SDN controller & Slicing Manager (CSC&SM). Particularly, FSC abstracts fog satellite nodes and manages their resources while CSC&SM optimizes resource management of network slices with a global view of the network. Moreover, a 5G Satellite Network Slice Management(5G SNSM) [68] is proposed to perform flexible network slice allocation based on QoS requirements. Simulation results show that the proposed architecture can improve the end-to-end delay and jitter by at least 7% and 50%, respectively.
@@ 3.1.3 Satellite Edge Computing
3.1.4 Slice Admission Control
Slice admission control is another important aspect of network slicing that should be considered simultaneously with resource allocation. In [79], a two-stage approach is developed to jointly optimize resource allocation and slice admission control in LEO satellite networks. To this end, the authors first formulate a robust optimization problem to optimize slice admission, taking into account the uncertainty in the network state and user distribution. Based on the admission decisions in the first stage, the resource allocation strategies are optimized in the second stage to guarantee QoS for users. Since the original optimization problem is non-convex, auxiliary variables are introduced to transform the problem into a convex form, thereby significantly reducing the problem’s complexity. Simulation results show that the proposed approach can improve the user QoS by 7.2% compared to the non-robust approach.
3.2 Integrated Satellite-Terrestrial Networks
@@ 3.2.1 Resource Allocation
Innovative approaches are proposed to overcome those challenges. To address the resource allocation problem in slice-aware ISTN, optimizing the VNE [82] algorithm is a promising solution. To that end, the authors in [81] design a framework, namely Slice-Aware VNE for Satellite-Terrestrial (SAST-VNE) to improve the implementation of the VNE algorithm in ISTNs. In this framework, problematic network slices, i.e., those that do not meet Key Performance Indicators (KPIs) or will undergo satellite handover, are inserted into a queue for analysis. In case a new slice is created, the embedding is computed using the DViNE algorithm [83]. For problematic network slices, if they have high priority, i.e., low tolerated latency, each affected link will be remapped with the shortest-path algorithm. On the contrary, if they have lower priority, the authors propose an iterative algorithm, namely SAST-VNE, to balance the network load and minimize the migration cost. Simulation results show that during handovers, SAST-VNE can reduce the average node migrations and average link migration by roughly 25% and 6%, respectively. Another implementation of VNE for ISTNs is presented in [80]. Particularly, the authors design a testbed to validate the feasibility of integrating non-GEO satellite constellations and implementing VNE algorithms in highly dynamic network conditions. This testbed consists of a dynamic satellite-terrestrial network emulated in Mininet [80], an external Ryu SDN [80] controller and a VNE algorithm script. First, the Mininet emulator is used to build an OpenFlow-based substrate network, enabling the Ryu SDN controller to manage the flow of packets in the network. The Ryu SDN controller also uses a Traffic Engineering (TE) application to create paths for each VN, set rate limits, gather network statistics, read the network topology, and process changes to the network topology in real-time. Furthermore, the VNE algorithm script, implemented in Matlab using an Integer Linear Programming formulation with a load-balancing objective function, handles required resources for virtual networks and updates the network configuration to the Ryu SDN controller. Simulation results show that the testbed can handle many scenarios and changes in the network topology such as adding new VNs, dynamic changes in the network layout, and handling of network failures.
@@
Moreover, approaches to optimize resource allocation are presented in [84] and [85]. Particularly, the authors in [84] propose a dynamic slicing strategy for ISTNs. The proposed strategy uses mirror nodes, instead of virtual nodes, to store motion duration and resource information of satellites. Based on that, the service forwarding path and slice resources can be adjusted, as illustrated in Fig. 10. The mirror nodes also communicate the satellite resources with ground stations and receive feedback, i.e., which satellites to be chosen for the slicing process. An optimization model is then formulated to maximize the long-term average slicing performance of the network with a constraint of computing, storage and bandwidth resources. Moreover, a resource reserved in slices and adaptive adjustment between slices algorithm (R2A2) is developed to solve the optimization problem. Simulation results show that, compared to a DQN approach, R2A2 can increase resource utilization and slice satisfaction by roughly 33% and 30%, respectively. With a focus on the admission decision for a slice request, the authors in [85] propose a game-theory-based solution for resource allocation in ISTNs. To solve the challenges related to limited resource constraints, slice admission control is formulated involving different combinations of service providers and users. Specifically, service providers and users form a multi-sided market to exchange resources. To participate in the network, both sides consider their prices, i.e., network resources for the providers and the cost for the users. An auction, designed with the Multi-Sided Ascending-Price Auction Mechanism [85], is performed by increasing the price for both sides based on supply and demand until a balance is reached. Eventually, the final prices are set, and the allocation of resources to users is determined. Experimental results show that, compared to a baseline approach, i.e., second-price auction [85], the ascending-price auction can increase the bandwidth per user, the admission ratio, and the gain from trade per user by up to 25%, 33%, and 20%, respectively.
@@
A novel approach for network slice orchestration is presented in [86]. Particularly, the authors present a system model for the deployment of network slicing for aircrafts using satellites. This model aims to enhance resource management in ISTNs by incorporating the concept of slice collaboration, i.e., defining network slices based on their willingness to share user traffic statistics with the infrastructure provider). To this end, a Mixed Integer Non Linear Programming (MINLP) model is developed to maximize the resources, i.e., cache and backhaul resources) while satisfying slice constraints. Moreover, a pricing model [86] is proposed to increase the chances of uncooperative slices, i.e., slices that are unwilling to share user traffic statistics, being served. Simulation results show that the MINLP model can increase the selection probability of uncooperative slices by 33% and cooperative slicing can accommodate 200% more services compared to that of uncooperative slicing. Taking another approach, the authors in [36] propose a slice-aware NTN architecture to enable the seamless integration between NTN and terrestrial networks. To this end, the authors propose an end-to-end slicing model which treats the NTN as a slice-aware link in the terrestrial network. Based on the proposed model, a functional architecture fully compliant with the 3GPP standard is developed to allow the NTN to be integrated with the terrestrial network. The architecture consists of three main segments: the user segment with the satellite terminal, the ground segment with satellite gateways and network control components, and the space segment with one or more satellites. The Satellite Terminal (ST) serves as the User Equipment (UE) in 5G networks, communicating with the space segment, which then relays messages to the satellite gateway. Satellite gateways are the central entities that manage resource allocation, authentication, and data processing. Simulation results show that the proposed slice-aware scheme can achieve better trip time, packet error rate, and jitter compared to those of the cases without network slicing.
3.2.2 User Assignment
@@
In addition to routing, several frameworks have been proposed to enhance network performance by optimizing user assignment in [87], [46], and [88]. Particularly, in [87], the authors propose a network slicing scheme for hybrid satellite-terrestrial networks, aiming to improve reliability and reduce the video traffic offload. To this end, the authors develop a scheduling strategy that assigns the users to three different types of networks, e.g., only satellite, only terrestrial, and hybrid satellite-terrestrial, based on Channel Quality Indicator (CQI) and QoS requirements. Simulation results show that the proposed scheduling strategy can improve the network throughput by up to 48% compared to the case where only the satellite network is employed. Differently, with a focus on space-terrestrial integrated vehicular networks (STIVN), the authors in [46] propose a scheme for network slicing to support both delay-tolerant services (DTSs) and delay-sensitive services (DSSs). This scheme aims to solve the problem of scheduling and resource slicing in STIVN, which involves allocating spectrum resources to slices, determining bandwidth allocation, and user assignment for each vehicle. Specifically, the system cost (i.e., DSS requirement violation, DTS delay, and slice reconfiguration) is taken into account in two subproblems: a resource slicing subproblem in large-timescale, and a resource scheduling subproblem in a smaller timescale. To solve those subproblems, a two-layered RL-based scheme is developed. In the first layer, namely resource slicing layer, spectrum-resources are pre-allocated to slices using a proximal policy optimization (PPO)-based RL algorithm [46]. In the second layer, namely resource scheduling layer, depending on network conditions and service requirements, spectrum resources are assigned to vehicles using match-based algorithms. Simulation results show that the proposed approach can reduce the overall system cost up to roughly 72.57%, compared to that of a baseline approach. In another approach, the authors in [88] propose a network slicing scheme to improve their existing resource management approach in satellite-LTE networks. In the previous approach, an adaptive hybrid satellite-LTE downlink scheduler (H-MUDoS) [89] determines if users can be served through satellite network or ground-based stations. However, simulation results show that the decrease in QoS of the satellite network affects the performance of the entire hybrid network. In the current approach, the network is separated into isolation slices, each with its own scheduling strategy assigned by the scheduler. Simulation results show that the decrease in QoS of a slice does not affect other slices.
3.2.3 Routing